# The initial stages of silver fluorination: a scanning tunneling microscopy investigation

### v1

@@ 3.1 Fluorination of Ag(100)
The apparent STM topography of isolated F adatoms on the surface strongly depends on the bias voltage, $V_{\rm B}$, and the tunneling current, $I_{\rm T}$. We first discuss the STM images collected with tunneling currents higher than 500 pA and later those taken with lower tunneling currents. Panels (a)-(h) in Fig. 2 show consecutive images of the same isolated F adatom measured with tunneling bias ranging from -1 to 1 V and a tunneling current of $I_{\rm T}=750$ pA. It is clear that the STM apparent shape of F adatoms is a sombrero when $V_{\rm B}<0.75$ V and becomes a depression at 1 V. The shape change is evident by looking at the STM pictures in Figs. 2 (a)-(h). At $V_{\rm B}=-1.0$ V, the F adatom presents the sombrero shape with a maximum in the apex which protrudes out from the Ag surface (withe region in the center of the topography in Fig. 2 (a)). It can be seen that, as $V_{\rm B}$ is increased, the height of the central protrusion of the sombrero is continuously reduced as well as the the depth of the depression around it, until the sombrero feature is lost, resulting in a smooth depression at 1.0 V, see Fig. 2 (h). The same evolution of the appearance change is observed in the sequence of STM images presented in B for a F adatom measured at different bias voltages after a fluorination process of 1.5 h.
@@ 3.2 Fluorination of Ag(110)
Figure 5: (a) Schematic representation of the first two surface layers of Ag(110). Large spheres represent surface atoms, while small spheres depict atoms in the subsurface layer. The red rectangle outlines the unit cell of the surface. Lattice parameters, denoted as $b$ and $b^{\prime}$, were determined from XRD patterns, reported in [41]. The four primary high-symmetry adsorption sites are indicated: hollow (hexagon), short bridge (square), long bridge (rhombus) and top (cross). (b) STM image showing some adsorbates in the top part and atomic resolution of the Ag(110) surface in the bottom part, acquired at $V_{\rm B}=-0.01$ V. The tunneling current was set to 0.1 and 1.0 nA in the top and bottom part, respectively. The green and pink dashed arrows indicate the directions along which the height profiles shown in panel (c) were taken. (d) Flattened topography of the same region shown in panel (b), optimized for better visualization. Red dots mark the positions of surface silver atoms, while white circle, black cross and cyan circle indicate the centers of a type-A, type-B and type-C adsorbates, respectively. The areas marked by the colored frames A, B and C are shown as three-dimensional renderings in the panels (e), (f) and (g), respectively. In each plot, a blue dot at the bottom highlights the position of the corresponding adsorbate.
@@
STM images, such as the one shown in Fig. 5, were analysed to pinpoint the adsorption sites of the adsorbates. In the top part of the topography, the three type of adsorbates are identified and marked: A (red circles), B (blue circles) and C (green circle). Since this image was acquired at $V_{\rm B}<0$, type-B adsorbates appear with a sombrero shape, while type-A and type-C adsorbates are observed as depressions with different depths. The deepest point for type-A and type-C adsorbates and the apex in the center of type-B adsorbates were determined from the right, left and center height profiles in Fig. 5 (c), respectively, and taken as the position of the adsorbates.
To determine the position of the Ag surface atoms in the top part of the image, the lattice observed in the bottom part was reproduced and marked with red dots in Fig. 5 (d). In this panel, the positions of the centers of type-A, type-B and type-C adsorbates, as extracted from the line-profiles in panel (c), are visible as white circles, a black cross and a cyan circle, respectively. This arrangement is better visualized in the three-dimensional representations displayed in Figs. 5 (e)-(g), and corresponding to the framed areas reported in panel (d). Here, the blue dot marks the position of each adsorbate, while the dashed line serves as a guide to eye, projecting this position onto the Ag(110) surface atomic plane (red dots). This shows that type-A adsorbates are located at the short-bridge sites on the Ag(110) surface, type-B at the hollow sites, whereas type-C at the top sites. A statistical analysis performed on several STM images reveals that the type-A adsorbates are the most abundant (60%), followed by type-B (35%). Type-C adsorbate is the least frequent, with only 5% of relative abundance.
@@ Appendix B STM images of isolated F adatom on Ag(100) as a function of the bias voltage
Typical STM images of F adatoms on Ag(100) are shown in Fig. 7. These topographies were recorded in the same area with bias voltages ranging from -2.0 to 1.0 V and a tunneling current of $I_{\rm T}=200$ pA. For $V_{\rm B}<0$, F adatoms appear as sombreros, following a similar transition from sombrero to depression as observed after a 3 h exposure, as shown in Figs. 2 (a)-(h) in 3.1.

### v2

@@ 3.1 Fluorination of Ag(100)
The apparent STM topography of isolated F adatoms on the surface strongly depends on the bias voltage, $V_{\rm B}$, and the tunneling current, $I_{\rm T}$. We first discuss the STM images collected with tunneling currents higher than 500 pA and later those taken with lower tunneling currents. Panels (a)-(h) in Fig. 2 show consecutive images of the same isolated F adatom measured with tunneling bias ranging from -1 to 1 V and a tunneling current of $I_{\rm T}=750$ pA. It is clear that the STM apparent shape of F adatoms is a sombrero when $V_{\rm B}<0.75$ V and becomes a depression at 1 V. The shape change is evident by looking at the STM pictures in Figs. 2 (a)-(h). At $V_{\rm B}=-1.0$ V, the F adatom presents the sombrero shape with a maximum in the apex which protrudes out from the Ag surface (white region in the center of the topography in Fig. 2 (a)). It can be seen that, as $V_{\rm B}$ is increased, the height of the central protrusion of the sombrero is continuously reduced as well as the the depth of the depression around it, until the sombrero feature is lost, resulting in a smooth depression at 1.0 V, see Fig. 2 (h). The same evolution of the appearance change is observed in the sequence of STM images presented in B for a F adatom measured at different bias voltages after a fluorination process of 1.5 h.
@@ 3.2 Fluorination of Ag(110)
Figure 5: (a) Schematic representation of the first two surface layers of Ag(110). Large spheres represent surface atoms, while small spheres depict atoms in the subsurface layer. The red rectangle outlines the unit cell of the surface. Lattice parameters, denoted as $b$ and $b^{\prime}$, were determined from XRD patterns, reported in [41]. The four primary high-symmetry adsorption sites are indicated: hollow (hexagon), short-bridge (square), long-bridge (rhombus) and top (cross). (b) STM image showing some adsorbates in the top part and atomic resolution of the Ag(110) surface in the bottom part, acquired at $V_{\rm B}=-0.01$ V. The tunneling current was set to 0.1 and 1.0 nA in the top and bottom part, respectively. The green and pink dashed arrows indicate the directions along which the height profiles shown in panel (c) were taken. (d) Flattened topography of the same region shown in panel (b), optimized for better visualization. Red dots mark the positions of surface silver atoms, while white circle, black cross and cyan circle indicate the centers of a type-A, type-B and type-C adsorbates, respectively. The areas marked by the colored frames A, B and C are shown as three-dimensional renderings in the panels (e), (f) and (g), respectively. In each plot, a blue dot at the bottom highlights the position of the corresponding adsorbate.
@@
STM images, such as the one shown in Fig. 5, were analysed to pinpoint the adsorption sites of the adsorbates. In the top part of the topography, the three type of adsorbates are identified and marked: A (red circles), B (blue circles) and C (green circle). Since this image was acquired at $V_{\rm B}<0$, type-B adsorbates appear with a sombrero shape, while type-A and type-C adsorbates are observed as depressions with different depths. The deepest point for type-A and type-C adsorbates and the apex in the center of type-B adsorbates were determined from the left, right and center height profiles in Fig. 5 (c), respectively, and taken as the position of the adsorbates.
To determine the position of the Ag surface atoms in the top part of the image, the lattice observed in the bottom part was reproduced and marked with red dots in Fig. 5 (d). In this panel, the positions of the centers of type-A, type-B and type-C adsorbates, as extracted from the line-profiles in panel (c), are visible as white dots, a black cross and a cyan dot, respectively. This arrangement is better visualized in the three-dimensional representations displayed in Figs. 5 (e)-(g), and corresponding to the framed areas reported in panel (d). Here, the blue dot marks the position of each adsorbate, while the dashed line serves as a guide to the eye, projecting this position onto the Ag(110) surface atomic plane (red dots). This shows that type-A adsorbates are located at the short-bridge sites on the Ag(110) surface, type-B at the hollow sites, whereas type-C at the top sites. A statistical analysis performed on several STM images reveals that the type-A adsorbates are the most abundant (60%), followed by type-B (35%). Type-C adsorbate is the least frequent, with only 5% of relative abundance.
@@ Appendix B STM images of isolated F adatom on Ag(100) as a function of the bias voltage
Typical STM images of F adatoms on Ag(100) are shown in Fig. 7. These topographies were recorded in the same area with bias voltages ranging from -2.0 to 1.0 V and a tunneling current of $I_{\rm T}=200$ pA. For $V_{\rm B}<0$, F adatoms appear as sombreros, following a transition from sombrero to depression as observed after a 3 h exposure, as shown in Figs. 2 (a)-(h) in 3.1.